Two trains are traveling on parallel tracks. Train A is traveling at 80 mph and train B is traveling at 88 mph. Train A passes the train station at 9:25. Train B passes the same station at 9:55. What time will it be when train B catches up with train A?
Answer by stanbon(75887)

Two trains are traveling on parallel tracks. Train A is traveling at 80 mph and train B is traveling at 88 mph. Train A passes the train station at 9:25. Train B passes the same station at 9:55. What time will it be when train B catches up with train A?
------------------
Train A DATA:
rate = 80 mph; time = x hrs ; distance = 80x miles
--------------------------------
Train B DATA:
rate = 88 mph; time = (x-(1/12)) hr. ; distance = 88(x-(1/12)) miles
------------------------------
EQUATION:
distance = distance
80x = 88(x-(1/12))
80x = 88x - (22/3)
8x = 22/3
x = 11/12 hr
x = 55 minutes
Ans: 9:25 + 55 minutes = 10:20
====================================
Cheers,
Stan H.